OpenStudy (anonymous):

What is the slope of -5x+2y=4 & -10x+4y=0? Are they parrallel, perpendicular or neither?

OpenStudy (anonymous):

Y= 5/2X +2 and Y= 5/2 x since the slope is the same it is parallel